OpenStudy (cassieforlife5):

Solve log4x + log4(x-2)= log4 15 Help please!

OpenStudy (sidsiddhartha):

use basic property of logarithms\[\log_{b} a+\log_{b} c=\log_{b} (ac)\]

OpenStudy (sidsiddhartha):

now can u use this on your problem?

OpenStudy (sidsiddhartha):

\[\log_{4} x+\log_{4} (x-2)=\log_{4} 15\] that is given right?

OpenStudy (sidsiddhartha):

here first two terms on the left hand side has a commom base of 4 ,so we can use the above property \[\log_{4} x(x-2)=\log_{4} 15\] ok?

OpenStudy (cassieforlife5):

Is the answer x= -3 or 5 ?

OpenStudy (sidsiddhartha):

answer is 5 because it will satisfy the equation

OpenStudy (sidsiddhartha):

\[x(x-2)=15\\if~X=5~u'\ll~have\\x(x-2)=5(5-2)=15\]

OpenStudy (sidsiddhartha):

ok got it?

OpenStudy (cassieforlife5):

Yeah. I didn't check them at the end. Thanks

OpenStudy (sidsiddhartha):

yes u have to check because when u'll get 2 answers 1 will not satisfy the main equation and i will satisfy